OpenStudy (anonymous):

A ball of mass 0.2 Kg rests on a vertical post of height 5m a bullet of mass 0.01 kg travelling with a Velcoity V hits the centre of ball in Horizontal after the collision the ball hit the ground at 20m and bullet at 100m find initial velocity of bullet?

OpenStudy (anonymous):

MV = m1u + Mu2

OpenStudy (anonymous):

assuming elastic collision call ball - M , bullet - m conservation of momentum : m*V = M * u + m * U

OpenStudy (anonymous):

Yup...i got it.....a fatal Mistake in my cal made me to post this question i am so silly..)

OpenStudy (anonymous):

now we know time to fall 5m is given by 5 = gt^2 / 2

OpenStudy (anonymous):

oh so you found the answer? :)